# Approximation Algorithms for Optimal Hopsets

### v1

@@ Abstract
A hopset $H$ with hopbound $\beta$ and stretch $\alpha$ for a given graph $G$ is a set of edges such that between every pair of vertices $u$ and $v$ there is a path with at most $\beta$ hops in $G\cup H$ that approximates the distance between $u$ and $v$ up to a multiplicative stretch of $\alpha$. Hopsets have found a wide range of applications in distance-based problems in various computational models since the 90s. More recently, there has been significant interest in understanding these fundamental objects from an existential and structural perspective. But all of this work takes a worst-case (or existential) point of view: How many edges do we need to add to satisfy a given hopbound and stretch requirement for any input graph?
We initiate the study of the natural optimization variant of this problem: given a specific graph instance, what is the minimum number of edges that satisfy the hopbound and stretch requirements? We give approximation algorithms for a generalized hopset problem which, when combined with known existential bounds, lead to different approximation guarantees for various regimes depending on hopbound, stretch, and directed vs. undirected inputs. We complement our upper bounds with a lower bound that implies Label Cover hardness for directed hopsets/shortcut sets with hopbound at least $3$.
@@ 1.1 Our Results and Techniques
Directed $\widetilde{O}(n^{2/5})$ Arbitrary $\widetilde{O}(\beta^{1/3}\cdot n^{2/3+\epsilon^{\prime}})$ 5.19
@@
Table 1: Comparison of hopset results, where $\epsilon^{\prime}>0$ is an arbitrarily small constant, $\epsilon\in(0,1)$, and $\eta>\beta^{1/(\ln{\ln{\beta}}-\frac{1}{2}\ln{\ln{\ln{\beta}}})}$. For $\beta\geq 3$, $\eta\geq 6$. “Arbitrary” stretch means arbitrary distance bounds, possibly different for each demand pair.
@@ 1.1.1 Upper Bounds
While this list may appear a bit complex, it turns out that almost all of the results are generated by trading off several approximation algorithms that we design and analyze with the known existential bounds for the given regime. In particular, we consider three main approximation algorithms: one based on rounding an LP relaxation, one based on randomly sampling stars, and one based on defining and analyzing an appropriate variant of junction trees [GKL23, GKL24, CDKL20]. Instead of just analyzing the approximation ratio of each algorithm as a function of $n$ (the traditional point of view), we analyze the approximation ratios as functions of $n$, $OPT$, $\beta$ and the local neighborhood size [DK11, BBM+11]. Importantly, none of these algorithms have performance that depends on whether the graph is undirected or directed, or what the allowed stretch is.
@@
Second, while the LP relaxation that we use for our rounding algorithm is a natural variant of the standard spanner LP relaxation (see [DK11, BBM+11, DNZ20]), the hopset variant turns out to be significantly more difficult to solve. Most notably, solving the equivalent LP relaxations for spanners (either the flow version of [DK11] or the fractional cut version of [DNZ20]) requires using the ellipsoid method with a separation oracle for a problem known as Restricted Shortest Path, in which we are given two distance metrics (think of one metric as being our original edge lengths and one metric as being the fractional values given by the LP to edges) and are asked to find the shortest path in the second metric subject to a distance constraint in the first metric. Unfortunately, Restricted Shortest Path is NP-hard, but it turns out that one can use approximation algorithms for Restricted Shortest Path to solve the LP up to any desired accuracy (see [DK11]). But for our hopset LP, the equivalent separation oracle has three metrics (the original edge lengths, the number of hops, and the LP values), and we need to find the shortest path in the third metric subject to upper bounds in the first two metrics. So we need to design a PTAS for this more complex problem in order to solve our LP.
Third, for similar reasons our junction tree argument is more complex. Junction trees were originally introduced for network design problems where the only constraints are on connectivity, not distances; see, e.g., [CEGS11]. By reducing to a “layered” graph, Chlamtáč et al. [CDKL20] showed that junction tree-based schemes are also useful in distance-constrained settings, and this layering technique was pushed significantly further by [GKL23]. When trying to use these ideas for hopsets, though, we have the same difficulty as when solving the LP: we have essentially two distance constraints for each demand, corresponding to the number of hops and the allowed distance. To overcome this, we give a “two-stage” layering reduction, where we first construct a layered graph that allows us to get rid of the hop requirement without changing the distance requirements. Then we can use the further layered graph of [GKL23] to get rid of the distance requirement, transforming it into a pure connectivity problem. In fact, we can use [GKL23] as a black-box for this second step.
@@ 1.1.2 Lower Bounds
Assuming that $NP\not\subseteq DTIME(2^{polylog(n)})$, for any constant $\epsilon>0$, and for any $\beta\geq 3$, there is no polynomial-time algorithm that can approximate directed Generalized $\beta$-Hopset (for any stretch value) or the minimum shortcut set on directed graphs with approximation ratio better than $2^{\log^{1-\epsilon}n}$.
@@ 1.2 Related Work
On the approximation algorithms side, most relevant to hopsets are the weighted pairwise spanner algorithms [GKL23], and 2-hop covers [CHKZ03]. In particular, weighted pairwise spanners (like hopsets and unlike $k$-spanners) do not have $n$ as a trivial lower bound on ${\sf OPT}$, and hence some of the difficulties encountered when designing algorithms are similar.
@@ 2 Preliminaries
Formally, we consider the following problem:
Definition 1 (Generalized $\beta$-Hopset).
Given a directed graph $G=(V,E)$, edge weights (or “lengths") $\ell:E\rightarrow\{1,2,3,...,\texttt{poly}(n)\}$, a set of vertex pairs $\mathcal{D}\subseteq V\times V$, and a distance bound function $Dist:\mathcal{D}\rightarrow\mathbb{N}_{\geq 0}$, find the smallest set of edges $H$ such that for each vertex pair $(s,t)\in\mathcal{D}$, it is the case that $d_{G}(s,t)\leq d^{(\beta)}_{H\cup G}(s,t)\leq Dist(s,t)$. In other words, there must be an $s-t$ path $P$ in $H\cup G$ such that $|P|\leq\beta$ and $\sum_{e\in P}\ell(e)\leq Dist(s,t)$.
${{\sf OPT}}$ to refer to the cost of the optimal solution to the Generalized $\beta$-Hopset problem instance. That is, OPT will refer to the number of edges in the optimal hopset. A path is an $i$-hop path if there are exactly $i$ edges on the path. We also say that a demand $(s,t)\in\mathcal{D}$ is settled or satisfied by a graph $G$ (or an edge set $E$) if there is a path from $s$ to $t$ in $G$ (in $E$) with at most $\beta$ hops and distance at most $Dist(s,t)$. Otherwise, demand $(s,t)$ is considered unsettled or unsatisfied.
@@ 3 LP Relaxation
Let $\mathcal{P}_{s,t}$ be the set of paths from $s$ to $t$ that have at most $\beta$ hops and path length at most $Dist(s,t)$. We will refer to these paths as “allowed” or “valid" paths. The natural flow LP for our problem requires building enough capacity to send one unit of (non-interfering) flow along valid paths for each demand; this is the basic LP used in essentially all network design problems, and was introduced for spanners by [DK11]. An equivalent LP, which is what we will use for Generalized $\beta$-Hopset, is obtained through the duality between flows and fractional cuts (of valid paths). This version of the LP for spanners was studied by [DNZ20], and strengthens the anti-spanner LP of [BBM+11].
@@ 3.1.2 Oracle 2: Hopbounded Restricted Shortest Path Problem
LP(Oracle 1) is also exponential in the number of constraints, so we use the ellipsoid algorithm again, with yet another separation oracle. Given an LP solution $\bm{\mathrm{z}}$, we must determine whether there is some valid $s-t$ path that is not fractionally cut (or, covered) by $\bm{\mathrm{z}}$. Specifically, we need to find a violated constraint of the form $\sum_{e\in P}z_{e}<1$ for some path $P\in\mathcal{P}_{s,t}$. Observe that we now have three separate metrics: the $\bm{\mathrm{z}}$ metric (given by a solution to LP(Oracle 1)), our original distance metric from the input, and a hop metric (where each edge has hop “length" 1). We only care about valid $s-t$ paths; namely, paths with length at most $D(s,t)$ in our original distance metric and with length at most $\beta$ in our hop metric. Thus, our goal is to find the shortest path in the $\bm{\mathrm{z}}$ metric that respects these upper bounds in the distance and hop metrics.
@@
The Restricted Shortest Path problem is well studied [Has92, War87, Phi93, XZTT08]. The problem admits an FPTAS, meaning there exists a polynomial-time (1+$\epsilon$)-approximation for the problem [LR01]. Since our problem of interest has a third metric (the hop metric), we refer to it as the Hopbounded Restricted Shortest Path problem. We will modify the Restricted Shortest Path FPTAS algorithm of [LR01] to give an FPTAS for the Hopbounded Restricted Shortest Path problem, thus giving a $(1+\epsilon)$-approximate separation oracle for LP(Oracle 1). We formally define the Hopbounded Restricted Shortest Path problem with respect to LP(Oracle 1):
@@ 3.1.3 Hopbounded Restricted Shortest Paths Algorithm
At a high level, the Restricted Shortest Path algorithm works as follows. The algorithm runs multiple binary searches to find good upper and lower bounds on the cost of the optimal solution, then uses these bounds to scale and discretize (or “bucket") the costs of the edges. They then give a pseudo-polynomial time dynamic programming algorithm on the problem with bucketed edge costs, which they show is a $(1+\epsilon)$-approximation for the original problem. For runtime, the pseudo-polynomial time DP algorithm is polynomial in $U/L$ (and in $1/\epsilon$), where $U$ and $L$ are the upper and lower bounds, respectively, on the optimal solution. The bounds they find are such that $U/L=O(n)$, and so the algorithm is an FPTAS.
@@ 4 Approximation for Hopbound 2
Fix some $(u,v)\in V\times V$. If $x_{(u,v)}\geq 1/(c\ln n)$ then our algorithm will directly include the $(u,v)$ edge. So without loss of generality, we may assume that $x_{(u,v)}\leq 1/2$. As discussed, our LP (LP(Hopset)) is equivalent to the flow LP, and hence $\bm{\mathrm{x}}$ supports flows $f:\mathcal{P}_{u,v}\rightarrow[0,1]$ such that $\sum_{P\in\mathcal{P}_{u,v}}f_{P}\geq 1$. Since $\beta=2$ and $x_{(u,v)}\leq 1/2$, this means that $\sum_{P\in\mathcal{P}_{u,v}\setminus\{(u,v)\}}f_{P}\geq 1/2$, i.e., at least $1/2$ flow is sent on paths of length $2$ in $\mathcal{P}_{u,v}$. Let $W\subseteq V\setminus\{u,v\}$ be the set of nodes that are the middle node of such a path, and for each $w\in W$ let $P_{w}$ denote the path $u-w-v\in\mathcal{P}_{u,v}$.
@@ 5 Approximation Algorithms for General Hopbounds
Continuing the connection to spanners, there is a reduction from Generalized $\beta$-Hopset to the directed pairwise weighted spanner problem (where “weighted" refers to edge costs). The most general version of this problem, studied by [GKL23], allows for any demand set, any positive rational edge costs, integer edge weights in $\texttt{poly}(n)$, and arbitrary distance bound functions. The reduction starts with the transitive closure $G_{M}$, and builds a layered graph with $\beta+1$ copies of each node in $G_{M}$, and $\beta$ copies of each edge (see Section 5.1.2 for a more detailed description of the reduction). Since [GKL23] achieved an $\widetilde{O}(n^{4/5+\epsilon})$-approximation for directed pairwise weighted spanners, this reduction immediately gives an $\widetilde{O}((\beta n)^{4/5+\epsilon})$-approximation for Generalized $\beta$-Hopset.
In this section, we improve upon this result and get an $\widetilde{O}(n^{4/5+\epsilon})$-approximation for Generalized $\beta$-Hopset, removing the dependence on $\beta$. We will give approximation algorithms in terms of $n,\beta,{{\sf OPT}}$, and “local neighborhood size." All of our algorithms are based on spanner algorithms, and we must modify them (and provide different analyses) to accommodate the hop constraint. We will then trade these algorithms off with known existential hopset results to get approximations (in terms of $n$ and $\beta$) in many regimes, including the general setting (directed graphs, arbitrary stretch), and in more restricted settings, such as small and large hopbound, and specific stretch regimes. The approximation we get in any given regime will be a function of how good the existential bounds are for that regime, so better existential results will result in better approximations.

### v2

@@ Abstract
For a given graph $G$, a hopset $H$ with hopbound $\beta$ and stretch $\alpha$ is a set of edges such that between every pair of vertices $u$ and $v$, there is a path with at most $\beta$ hops in $G\cup H$ that approximates the distance between $u$ and $v$ up to a multiplicative stretch of $\alpha$. Hopsets have found a wide range of applications for distance-based problems in various computational models since the 90s. More recently, there has been significant interest in understanding these fundamental objects from an existential and structural perspective. But all of this work takes a worst-case (or existential) point of view: How many edges do we need to add to satisfy a given hopbound and stretch requirement for any input graph?
We initiate the study of the natural optimization variant of this problem: given a specific graph instance, what is the minimum number of edges that satisfy the hopbound and stretch requirements? We give approximation algorithms for a generalized hopset problem which, when combined with known existential bounds, lead to different approximation guarantees for various regimes depending on hopbound, stretch, and directed vs. undirected inputs. We complement our upper bounds with a lower bound that implies Label Cover hardness for directed hopsets and shortcut sets with hopbound at least $3$.
@@ 1.1 Our Results and Techniques
Directed $\widetilde{O}(n^{2/5})$ Individual $\widetilde{O}(\beta^{1/3}\cdot n^{2/3+\epsilon^{\prime}})$ 5.19
@@
Table 1: Comparison of hopset results, where $\epsilon^{\prime}>0$ is an arbitrarily small constant, $\epsilon\in(0,1)$, and $\eta>\beta^{1/(\ln{\ln{\beta}}-\frac{1}{2}\ln{\ln{\ln{\beta}}})}$. For $\beta\geq 3$, $\eta\geq 6$. “Individual” stretch means arbitrary distance bounds, possibly different for each individual demand pair. All results are for pairwise demands, with edge lengths being nonnegative integer and upper bounded by $\texttt{poly}(n)$.
Demands Edge Lengths Edge Costs? Stretch Approximation Paper
Pairwise Arbitrary Y 2 $O(\ln n)$ [DK11]
All-Pairs Arbitrary N Arbitrary $\widetilde{O}(n^{1/2})$ [BBM+11]
Pairwise Integer, $\leq\texttt{poly}(n)$ Y Individual $\widetilde{O}(n^{4/5+\epsilon})$ [GKL23]
Pairwise Vector Y Indiv. Vectors $\widetilde{O}(|D|^{1/2+\epsilon}\cdot\tau^{m-1})$ [GKL24b]
Table 2: Comparison of related results for directed spanners. The result of [GKL24b] allows for edge resource consumption vectors as opposed to edge lengths. $D$ is the set of demands, $\epsilon>0$ is an arbitrarily small constant, $\tau$ is the largest resource budget, and $m$ is the size of the resource consumption vectors. “Individual” stretch means arbitrary distance bounds, possibly different for each individual demand pair. All variants require nonnegative edge lengths.
@@ 1.1.1 Upper Bounds
While this list may appear a bit complex, it turns out that almost all of the results are generated by trading off several approximation algorithms that we design and analyze with the known existential bounds for the given regime. In particular, we consider three main approximation algorithms: one based on rounding an LP relaxation, one based on randomly sampling stars, and one based on defining and analyzing an appropriate variant of junction trees [GKL23, GKL24b, CDKL20]. Instead of just analyzing the approximation ratio of each algorithm as a function of $n$ (the traditional point of view), we analyze the approximation ratios as functions of $n$, $OPT$, $\beta$ and the local neighborhood size [DK11, BBM+11]. Importantly, none of these algorithms have performance that depends on whether the graph is undirected or directed, or what the allowed stretch is.
@@
We also note that there is a simple reduction from hopsets to the multicriteria spanner problem, introduced by [GKL24b]. For multicriteria spanners, instead of edges being associated with just edge lengths—as with traditional spanners—edges in the multicriteria spanner problem each have a resource consumption vector, where each entry corresponds to the consumption of the respective resource on that edge. The simple reduction to multicriteria spanners implies a $\widetilde{O}(|D|^{1/2+\epsilon}\cdot\beta)$-approximation for hopsets, where $D$ is the set of vertex demand pairs given in the input (see Table 2).
@@
Second, while the LP relaxation that we use for our rounding algorithm is a natural variant of the standard spanner LP relaxation (see [DK11, BBM+11, DNZ20]), the hopset variant turns out to be significantly more difficult to solve. Most notably, solving the equivalent LP relaxations for spanners (either the flow version of [DK11] or the fractional cut version of [DNZ20]) requires using the ellipsoid method with a separation oracle for a problem known as Restricted Shortest Path, in which we are given two distance metrics (think of one metric as being our original edge lengths and one metric as being the fractional values given by the LP to edges) and are asked to find the shortest path in the second metric subject to a distance constraint in the first metric. Unfortunately, Restricted Shortest Path is NP-hard, but it turns out that one can use approximation algorithms for Restricted Shortest Path to solve the LP up to any desired accuracy (see [DK11]). But for our hopset LP, the equivalent separation oracle has three metrics (the original edge lengths, the number of hops, and the LP values), and we need to find the shortest path in the third metric subject to upper bounds in the first two metrics. So we design a PTAS for this more complex problem in order to solve our LP.
Third, for similar reasons our junction tree argument is more complex. Junction trees were originally introduced for network design problems where the only constraints are on connectivity, not distances; see, e.g., [CEGS11, FKN12]. By reducing to a “layered” graph, Chlamtáč et al. [CDKL20] showed that junction tree-based schemes are also useful in distance-constrained settings, and this layering technique was pushed significantly further by [GKL23]. When trying to use these ideas for hopsets, though, we have the same difficulty as when solving the LP: we have essentially two distance constraints for each demand, corresponding to the number of hops and the allowed distance. To overcome this, we give a “two-stage” layering reduction, where we first construct a layered graph that allows us to get rid of the hop requirement without changing the distance requirements. Then we can use the further layered graph of [GKL23] to get rid of the distance requirement, transforming it into a pure connectivity problem. In fact, we can use [GKL23] as a black-box for this second step.
@@ 1.1.2 Lower Bounds
Assuming that $NP\not\subseteq DTIME(2^{polylog(n)})$, for any constant $\epsilon>0$, and for any $\beta\geq 3$, there is no polynomial-time algorithm that can approximate directed Generalized $\beta$-Hopset (for any stretch value; see Definition 2) or the minimum shortcut set on directed graphs with approximation ratio better than $2^{\log^{1-\epsilon}n}$.
@@ 1.2 Related Work
On the approximation algorithms side, most relevant to hopsets are the weighted pairwise spanner algorithms [GKL23] (see Table 2), 2-hop covers [CHKZ03], and multicriteria spanners [GKL24b] (Table 2). In particular, weighted pairwise spanners (like hopsets and unlike $k$-spanners) do not have $n$ as a trivial lower bound on ${\sf OPT}$, and hence some of the difficulties encountered when designing algorithms are similar. Additionally, when we do not have distance bounds—that is, when stretch is $\infty$ for all demand pairs—weighted pairwise spanners captures the hopset problem.
@@
Multicriteria spanners were recently introduced by [GKL24b]; they also rely on a modification of the junction tree framework to give an approximation algorithm. While their junction tree framework works for our setting, we present a different framework tailored to hopsets that yields better approximations for our setting than their framework gives. Other related problems are optimizations for other variants of spanners, including traditional directed $k$-spanner (unit edge costs, all-pairs demands) [DK11, BBM+11] (Table 2), buy-at-bulk spanners [GKL24a], and transitive closure spanners [BGJ+08, CJMN25]. In addition to transitive closure spanners, [CJMN25] shows hardness of bicriteria approximation for shortcut sets (where they allow the hopbound to be violated).
@@ 2 Preliminaries
Formally, we consider the following problem:
Definition 2 (Generalized $\beta$-Hopset).
Given a directed graph $G=(V,E)$, edge weights (or “lengths”) $\ell:E\rightarrow\{1,2,3,...,\texttt{poly}(n)\}$, a set of vertex pairs $\mathcal{D}\subseteq V\times V$, and a distance bound function $Dist:\mathcal{D}\rightarrow\mathbb{N}_{\geq 0}$, find the smallest set of edges $H\subseteq\widetilde{E}$ such that for each vertex pair $(s,t)\in\mathcal{D}$, it is the case that $d_{G}(s,t)\leq d^{(\beta)}_{H\cup G}(s,t)\leq Dist(s,t)$. In other words, there must be an $s-t$ path $P$ in $H\cup G$ such that $|P|\leq\beta$ and $\sum_{e\in P}\ell(e)\leq Dist(s,t)$.
We will use ${{\sf OPT}}$ to refer to the cost of the optimal solution to the Generalized $\beta$-Hopset problem instance. That is, OPT will refer to the number of edges in the optimal hopset. A path is an $i$-hop path if there are exactly $i$ edges on the path. We also say that a demand $(s,t)\in\mathcal{D}$ is settled or satisfied by a graph $G$ (or an edge set $E$) if there is a path from $s$ to $t$ in $G$ (in $E$) with at most $\beta$ hops and distance at most $Dist(s,t)$. Otherwise, demand $(s,t)$ is considered unsettled or unsatisfied.
@@ 3 LP Relaxation
Let $\mathcal{P}_{s,t}$ be the set of paths from $s$ to $t$ that have at most $\beta$ hops and path length at most $Dist(s,t)$. We will refer to these paths as “allowed” or “valid” paths. The natural flow LP for our problem requires building enough capacity to send one unit of (non-interfering) flow along valid paths for each demand; this is the basic LP used in essentially all network design problems, and was introduced for spanners by [DK11]. An equivalent LP, which is what we will use for Generalized $\beta$-Hopset, is obtained through the duality between flows and fractional cuts (of valid paths). This version of the LP for spanners was studied by [DNZ20], and strengthens the anti-spanner LP of [BBM+11].
@@ 3.1.2 Oracle 2: Hopbounded Restricted Shortest Path Problem
LP(Oracle 1) is also exponential in the number of constraints, so we use the ellipsoid algorithm again, with yet another separation oracle. Given an LP solution $\bm{\mathrm{z}}$, we must determine whether there is some valid $s-t$ path that is not fractionally cut (or, covered) by $\bm{\mathrm{z}}$. Specifically, we need to find a violated constraint of the form $\sum_{e\in P}z_{e}<1$ for some path $P\in\mathcal{P}_{s,t}$. Observe that we now have three separate metrics: the $\bm{\mathrm{z}}$ metric (given by a solution to LP(Oracle 1)), our original distance metric from the input, and a hop metric (where each edge has hop “length” 1). We only care about valid $s-t$ paths; namely, paths with length at most $D(s,t)$ in our original distance metric and with length at most $\beta$ in our hop metric. Thus, our goal is to find the shortest path in the $\bm{\mathrm{z}}$ metric that respects these upper bounds in the distance and hop metrics.
@@
The Restricted Shortest Path problem is well studied [Has92, War87, Phi93, XZTT08, HK18]. The problem admits an FPTAS, meaning there exists a polynomial-time (1+$\epsilon$)-approximation for the problem [LR01]. Since our problem of interest has a third metric (the hop metric), we refer to it as the Hopbounded Restricted Shortest Path problem. We will modify the Restricted Shortest Path FPTAS algorithm of [LR01] to give an FPTAS for the Hopbounded Restricted Shortest Path problem, thus giving a $(1+\epsilon)$-approximate separation oracle for LP(Oracle 1). We formally define the Hopbounded Restricted Shortest Path problem with respect to LP(Oracle 1):
@@ 3.1.3 Hopbounded Restricted Shortest Paths Algorithm
At a high level, the Restricted Shortest Path algorithm works as follows. The algorithm runs multiple binary searches to find good upper and lower bounds on the cost of the optimal solution, then uses these bounds to scale and discretize (or “bucket”) the costs of the edges. They then give a pseudo-polynomial time dynamic programming algorithm on the problem with bucketed edge costs, which they show is a $(1+\epsilon)$-approximation for the original problem. For runtime, the pseudo-polynomial time DP algorithm is polynomial in $U/L$ (and in $1/\epsilon$), where $U$ and $L$ are the upper and lower bounds, respectively, on the optimal solution. The bounds they find are such that $U/L=O(n)$, and so the algorithm is an FPTAS.
@@ 4 Approximation for Hopbound 2
Fix some $(u,v)\in V\times V$. If $x_{(u,v)}\geq 1/(c\ln n)$ then our algorithm will directly include the $(u,v)$ edge. So without loss of generality, we may assume that $x_{(u,v)}\leq 1/2$. As discussed, (LP(Hopset)) is equivalent to the flow LP, and hence $\bm{\mathrm{x}}$ supports flows $f:\mathcal{P}_{u,v}\rightarrow[0,1]$ such that $\sum_{P\in\mathcal{P}_{u,v}}f_{P}\geq 1$. Since $\beta=2$ and $x_{(u,v)}\leq 1/2$, this means that $\sum_{P\in\mathcal{P}_{u,v}\setminus\{(u,v)\}}f_{P}\geq 1/2$, i.e., at least $1/2$ flow is sent on paths of length $2$ in $\mathcal{P}_{u,v}$. Let $W\subseteq V\setminus\{u,v\}$ be the set of nodes that are the middle node of such a path, and for each $w\in W$ let $P_{w}$ denote the path $u-w-v\in\mathcal{P}_{u,v}$.
@@ 5 Approximation Algorithms for General Hopbounds
Continuing the connection to spanners, there is a reduction from Generalized $\beta$-Hopset to the directed pairwise weighted spanner problem (where “weighted” refers to edge costs). The most general version of this problem, studied by [GKL23], allows for any demand set, any positive rational edge costs, integer edge weights in $\texttt{poly}(n)$, and arbitrary distance bound functions. The reduction starts with the transitive closure $G_{M}$, and builds a layered graph with $\beta+1$ copies of each node in $G_{M}$, and $\beta$ copies of each edge (see Section 5.1.2 for a more detailed description of the reduction). Since [GKL23] achieved an $\widetilde{O}(n^{4/5+\epsilon})$-approximation for directed pairwise weighted spanners, this reduction immediately gives an $\widetilde{O}((\beta n)^{4/5+\epsilon})$-approximation for Generalized $\beta$-Hopset.
In this section, we improve upon this result and get an $\widetilde{O}(n^{4/5+\epsilon})$-approximation for Generalized $\beta$-Hopset, removing the dependence on $\beta$. We will give approximation algorithms in terms of $n,\beta,{{\sf OPT}}$, and “local neighborhood size.” All of our algorithms are based on spanner algorithms, and we must modify them (and provide different analyses) to accommodate the hop constraint. We will then trade these algorithms off with known existential hopset results to get approximations (in terms of $n$ and $\beta$) in many regimes, including the general setting (directed graphs, arbitrary stretch), and in more restricted settings, such as small and large hopbound, and specific stretch regimes. The approximation we get in any given regime will be a function of how good the existential bounds are for that regime, so better existential results will result in better approximations.